OpenStudy (anonymous):

Find the zeros of the function f(x) = x2 - 5x - 6 by factoring. What is the largest zero of the function? Enter a number only.

hero (hero):

Hint: Let f(x) = 0

OpenStudy (anonymous):

??

OpenStudy (anonymous):

0=x2-5x-6

hero (hero):

Yes, correct

OpenStudy (anonymous):

and then what?

hero (hero):

2nd Hint: Split the middle term 0 = x^2 - 6x + 1x - 6

hero (hero):

Are you familiar with factor by grouping?

OpenStudy (anonymous):

can u solve it 4 me n show me how you got ur awnser?

OpenStudy (callisto):

x^2 - 6x + 1x - 6 = x^2 +1x -6x - 6 = x(x+1) -6(x+1) Can you try the next step?

OpenStudy (anonymous):

im not sure wat 2 do

OpenStudy (callisto):

Take out the common factor

OpenStudy (lgbasallote):

wow...this method's new to me..thanks @Callisto and @Hero ^_^ i learned something new hahaha